[Hallc_running] [New Comment] Hall C status update (Dec. 7)

gaskelld at jlab.org gaskelld at jlab.org
Thu Dec 7 16:05:01 EST 2017

 - by gaskelld on Thu, 12/07/2017 - 16:04.

Update @ 4pm:  - Update @ 4pm: 

The 5th pass separator issue appears to have been addressed.  After recovery from the earlier CHL crash, the program for tonight will be beam restoration to Hall D then 5th pass setup for Hall B.
Setup for A and C scheduled for day shift Friday. Shifts are still as noted earlier (target operator for Thursday wing and Friday owl - all shift workers should report Friday Day).


Logentry Text:
We are still recovering from the ESR trip yesterday afternoon - the HMS and SHMS magnets are on cold return, but still recovering.  The cryotarget cooldown started earlier this morning and is ongoing - expected completion by noon. At the moment we are in Controlled Access in support of the target cooldown.

Last night, some issues with the 5th pass separator resulted in some beam setup for Hall C. This was halted due in part to some issues with an unresolved FSD fault.  This latter issue will be addressed today.  

Once the target cooldown is complete, we will go to Beam Permit - depnding on the plan of action with regards to the 5th pass separator, beam setup to Hall C may resume.

For now, I would like to keep SL and 3rd shift workers "on call" until Friday, Day shift.  Target operators are required for all shifts now that the target is cooling down again. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3500722
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20171207/d2227523/attachment.html>

More information about the Hallc_running mailing list